OpenStudy (anonymous):

y=-2x+6,3y-x+3=0....Can somebody please explain how to solve this system using substitution???

OpenStudy (whpalmer4):

Since one of your equations directly gives you y in terms of x, replace y in the second equation with its equivalent in terms of x. Now solve for x. When you've got a value for x, plug it into the first equation to get y. It's really quite easy :-)

OpenStudy (whpalmer4):

Hint: \[3(-2x+6)-x+3=0\]

OpenStudy (anonymous):

Y=-2x+6 3y-x+3=0 Substitute in the other equation for y 3(-2x+6)-x+3=0 now multiply the parenthesis -6x+18-x+3=0 Now combine like terms -7x+21=0 add -7x to both sides 21=-7x divide both sides by -7x 3=x now figure Y so.. Y=-2(3)+6 y=-6+6 y=0 Checking: 3y - x + 3 = 0 3(0) - 3 + 3 = 0 0-3+3=0

OpenStudy (anonymous):

Thank you both :)